# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

### v1

@@ 4 For Engineering: Factors Influencing Carbon Credit Prices
In contrast, China’s carbon market is influenced by more policy and structural factors. Carbon regulation intensity is a critical factor in China’s carbon market, as the implementation of carbon regulation policies has a direct short-term impact on carbon prices, particularly by increasing compliance costs for organisations and suppressing market demand [35]. Crude oil and coal prices also play a crucial role in China’s market, with coal price fluctuations being especially prominent, reflecting China’s reliance on coal in its energy structure [44]. In addition, exchange rates and interest rates significantly impact China’s carbon market, particularly in the short and medium term, as fluctuations in exchange rates and interest rates affect market liquidity and energy prices [44, 35]. Lastly, the growth of industrial production and renewable energy also has a significant influence on China’s carbon prices. These factors reflect the level of industrialization in China and its ongoing energy transition, with increases in industrial production and renewable energy generation being closely linked to carbon market prices [25, 40].
@@ 5 Algorithms for Predicting Carbon Credit Prices
After thoroughly reviewing the factors influencing carbon prices, it becomes clear that these factors are not only critical for understanding carbon market fluctuations but also serve as essential input variables for carbon price prediction algorithms. Effectively incorporating these factors into predictive models can significantly enhance the accuracy of carbon price forecasts, making the review of these algorithms even more meaningful and necessary. By accurately predicting carbon prices, organisations can strategically plan their carbon credit purchases when prices are lower, thereby better managing their costs. This not only helps organisations control overall expenditures and strengthen their competitiveness in the market but also encourages them to actively participate in carbon management and transparently disclose their carbon emissions. In doing so, organisations can gain greater trust and advantage in markets that increasingly prioritize ESG principles. The following sections will now review the latest and effective carbon price prediction algorithms.
@@
Sun and Wang used a combined method of empirical mode decomposition [49] and least squares support vector machine to predict carbon prices. They proposed a hybrid model based on factor analysis, empirical mode decomposition, improved particle swarm optimization [50], and least squares support vector machine [51] to consider historical carbon prices and external factors influencing carbon prices. The study investigated three typical carbon markets in China, and the results showed that this hybrid model was more accurate in predicting carbon prices than other comparable models [52].
@@
Nadirgil employed the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) [53] technique to decompose the carbon price time series into multiple Intrinsic Mode Functions (IMFs), capturing various frequency characteristics within the data. Subsequently, Permutation Entropy (PE) was utilized to classify the IMFs, grouping IMFs with similar complexity into Cooperative Intrinsic Mode Functions (Co-IMFs) to reduce model computational complexity and the risk of overfitting. For high-frequency Co-IMFs, Variational Mode Decomposition (VMD) was further applied for secondary decomposition to mitigate the impact of high-frequency noise on model prediction. Ultimately, the study integrated multiple machine learning models, including Long Short-Term Memory (LSTM), Gated Recurrent Unit (GRU), Multilayer Perceptron (MLP), and Backpropagation Neural Network (BPNN), and optimized model parameters using Genetic Algorithm (GA) to construct 48 different hybrid prediction models. The effectiveness of the dual decomposition method (CEEMDAN-VMD-BPNN-GA) in addressing prediction complexity was ultimately demonstrated [54].
@@
Hao and Tian proposed an innovative carbon price prediction framework. First, the framework applies ICEEMDAN [55] to decompose the carbon price time series, effectively removing noise while retaining key signal features. Then, feature selection is conducted using the Partial Autocorrelation Function (PACF) and Maximum Relevance Minimum Redundancy (mRMR) methods. Next, the parameters of the KELM model are optimized using the Multi-Objective Chaotic Sine-Cosine Algorithm (MOCSCA). Finally, a multi-step prediction strategy is designed using the recursive method, which not only accurately predicts short-term price fluctuations but also provides reliable support for long-term trend forecasting [56].
@@
Li and Liu proposed a carbon price prediction model based on secondary decomposition and feature selection. First, the carbon price time series is subjected to multi-level decomposition using Discrete Wavelet Transform (DWT) [57] and ICEEMDAN, breaking it down into sub-sequences of different frequencies. Simultaneously, a three-stage feature selection method is employed to identify external influencing factors that are highly correlated with carbon prices. Subsequently, Support Vector Regression (SVR) [58] and Multilayer Perceptron (MLP) are used to predict the high-frequency and low-frequency sub-sequences, respectively, in order to capture short-term fluctuations and long-term trends. Finally, a weighted ensemble of the predictions from each sub-sequence is performed to obtain the overall carbon price prediction [59].
@@
Wang et al. proposed a comprehensive carbon price prediction framework. Initially, the carbon price time series undergoes multi-level decomposition and denoising using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) and Wavelet Transform (WT) to extract key features and reduce noise interference. Subsequently, Principal Component Analysis (PCA), Random Forest (RF), and Gradient Boosting Decision Tree (GBDT) are employed for feature selection, ensuring that the variables input into the model possess the strongest predictive capabilities. In the prediction phase, Bidirectional Gated Recurrent Unit (BIGRU) [60] is utilized for point prediction, leveraging bidirectional layers to capture the complex dynamic patterns within the time series. Following this, Gaussian Process Regression (GPR) [61] is applied for interval prediction, providing confidence intervals for the forecast results [62].
@@
Liu and Shen proposed a carbon price prediction model based on Empirical Wavelet Transform (EWT) [63] and Gated Recurrent Unit (GRU), aiming to enhance prediction accuracy by integrating multiple techniques. First, the carbon price time series is decomposed using EWT, extracting components of different frequencies, including trend components, low-frequency components, and high-frequency components. Subsequently, Fuzzy C-Means (FCM) [64] clustering is employed to classify these components, further refining the signal characteristics. Then, the Partial Autocorrelation Function (PACF) is used to analyze the lag relationships of each component, determining the optimal lag order as input for the GRU model. Finally, the GRU model performs single-step predictions for each component, and the predicted results are aggregated to obtain the overall carbon price prediction [65].
@@
Huang and He proposed a carbon price prediction model based on multi-scale joint prediction, integrating techniques such as Mean Optimization Empirical Mode Decomposition (MOEMD), Grey Relational Analysis (GRA), Factor Analysis (FA), Cooperative Kidney Algorithm (CKA) [66], and Extreme Learning Machine (ELM). First, MOEMD is used to decompose the carbon price data, extracting modal components at different frequencies. Then, GRA and FA are employed to select both structured and unstructured data highly correlated with carbon price fluctuations, including unstructured data like Baidu Search Index. Subsequently, CKA is utilized to optimize the parameters of the ELM model, enhancing prediction accuracy and stability. The experimental results demonstrate that the proposed MOEMD-CKA-ELM model significantly outperforms other comparative models in carbon price prediction [67].
@@
Yang et al. proposed a carbon price prediction model based on a multi-factor combination, integrating Modified Ensemble Empirical Mode Decomposition (MEEMD), Long Short-Term Memory (LSTM) networks [68], and a rule-based machine reasoning system. First, MEEMD is employed to decompose the carbon price time series, extracting modal components of different frequencies, which provide clear signals for subsequent predictions [69]. Then, the LSTM model, with its strong capability for processing time series data, accurately predicts each modal component. To further enhance prediction accuracy, a rule-based machine reasoning system is introduced to automatically optimize the key parameters of the LSTM model, ensuring robust performance under various conditions. The experimental results demonstrate that the proposed MEEMD-LSTM model outperforms traditional models in carbon price prediction [70].
@@
Lu et al. proposed a carbon price and trading volume prediction framework based on multi-model comparison, incorporating the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) technique to evaluate the performance of six different machine learning models. The machine learning models used include XGBoost [71], Random Forest (RF) [72], Grey Wolf Optimizer-based Nonlinear Extension of the Arps Decline Model (GWO-KNEA) [73], Particle Swarm Optimization-based Support Vector Machine (PSO-SVM) [74], Simulated Annealing and Fruit Fly Optimization Algorithm-based Support Vector Machine (SA-FFOA-SVM), and Radial Basis Function Neural Network (RBFNN) [75]. Through experiments conducted on data from eight carbon trading markets in China, the authors compared the prediction accuracy and stability of each model. The study found that the CEEMDAN-GWO-KNEA and CEEMDAN-RBFNN models performed exceptionally well across multiple datasets [76].
@@
Tian et al. proposed a novel hybrid model based on feature selection and multi-objective optimization algorithms for carbon price prediction. The study used Weighted Regularized Extreme Learning Machine (WRELM) as the artificial intelligence algorithm to improve the accuracy and stability of carbon price prediction [77]. The model overcame the limitations of statistical models in handling carbon price prediction and optimized the model using the Multi-Objective Grasshopper Optimization Algorithm (MOGOA), achieving excellent prediction results [78]. This research provides important references for the rational and stable development of the carbon market [79].
@@
Yang et al. proposed a carbon price prediction model based on Modified Ensemble Empirical Mode Decomposition (MEEMD), Long Short-Term Memory (LSTM) networks, and an Improved Whale Optimization Algorithm (IWOA) [80]. First, MEEMD is used to decompose the carbon price time series into several Intrinsic Mode Functions (IMFs) and a residual component, reducing noise and mode mixing issues [81]. Subsequently, the LSTM model is employed to predict the future values of each IMF and the residual component, capturing long-term dependencies in the time series [68]. To further enhance prediction performance, IWOA is used to optimize the initial weights and biases of the LSTM model [82].
@@
Zhao et al. proposed a carbon price prediction model that combines the Hodrick-Prescott (HP) filter, an improved Grey Model (Variable Grey Model, VGM), and Extreme Learning Machine (ELM) [83]. First, the HP filter is used to decompose the carbon price data and its influencing factors into long-term trends and short-term fluctuations, which are modeled separately. The VGM model is then employed to predict the long-term trends, while the ELM handles the short-term fluctuations. By combining the predictions from these two components, the model effectively enhances the accuracy of carbon price prediction [84].
@@
Wang et al. proposed a carbon price interval prediction model based on Lower Upper Bound Estimation (LUBE), Asymmetric Multi-Objective Evolutionary Algorithm (MOEA), and Long Short-Term Memory (LSTM) networks [68]. First, the carbon price time series is decomposed using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) to reduce noise and extract key features. Next, Symbol Transfer Entropy (STE) is employed for causal inference and feature selection to ensure the effectiveness of the input variables [85]. Then, based on the prediction points generated by LSTM, the LUBE model uses MOEA to optimize parameters and generate asymmetric prediction intervals [86].
@@
Deng et al. proposed a carbon price trend prediction model that integrates several advanced techniques. The model employs Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) to decompose carbon price time series into Intrinsic Mode Functions (IMFs) and residuals, allowing the extraction of significant features. Following this, a Two-Stage Feature Selection (TFS) method is used to identify the most relevant features from these IMFs, which are then fed into a Light Gradient Boosting Machine (LightGBM) model [87]. The hyperparameters of the LightGBM model are optimized using Bayesian Optimization Algorithm (BOA) to enhance its predictive performance [88]. Finally, SHapley Additive exPlanations (SHAP) is applied to interpret the model’s predictions [89], providing insights into the contributions of various features to the predicted carbon price trends [90].
@@
Shi et al. proposed a deep learning framework combining Convolutional Neural Networks (CNN) and Long Short-Term Memory (LSTM) networks to predict carbon prices with high accuracy and robustness. The model first uses CNN layers to extract spatial features from the carbon price time series data, which are then passed to LSTM layers to capture temporal dependencies [91]. This hybrid CNN-LSTM model is designed to address the limitations of traditional models in capturing complex nonlinear patterns in carbon markets [92]. The study demonstrates that the CNN-LSTM model significantly outperforms single CNN or LSTM models in both prediction accuracy and robustness, particularly in handling anomalies and outliers in the data [93].
@@
Zhang et al. proposed a hybrid model for multi-step carbon price prediction, integrating Extremely Randomized Trees (ET), Multivariate Variational Mode Decomposition (MVMD), and Long Short-Term Memory (LSTM) networks. The model employs ET to select the most influential external factors affecting carbon price fluctuations, simplifying the model’s complexity and enhancing prediction accuracy. [94] MVMD is then used to decompose the carbon price and selected external variables into multiple modes, allowing for the extraction of stable frequency components that are crucial for accurate predictions. Each mode is subsequently modeled using LSTM to capture long-term dependencies in the time series, with the final prediction obtained by aggregating the individual predictions [95].
@@
Zhang et al. proposed an enhanced decomposition-ensemble model for both point and probabilistic carbon price forecasting. The model integrates Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN), Fuzzy Entropy (FE) [96], Variational Mode Decomposition (VMD) [97], Long Short-Term Memory (LSTM) networks, Gated Recurrent Unit (GRU) networks, Cuckoo Search Algorithm (CS) [98], and Gaussian Process Regression (GPR). Initially, CEEMDAN is employed to decompose the carbon price series into Intrinsic Mode Functions (IMFs) and residuals, followed by FE to quantify the complexity of these IMFs. High-complexity IMFs undergo further decomposition using VMD. LSTM and GRU models then predict each decomposed component, with CS optimizing their weights. Finally, GPR is used to generate prediction intervals, offering probabilistic insights [99].
@@
Hong et al. proposed a hybrid carbon price forecasting framework that integrates a Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer) with a Multimodel Optimization Piecewise Error Correction System [100]. The DA-FEDformer model enhances frequency domain information extraction using a multi-layer perceptron integrated into the encoder-decoder layers, combined with an Improved Kernel Mean Square Error (IKMSE) loss function and the Lion Optimizer. This model generates initial carbon price predictions, which are then refined by the error correction system [55]. The system employs Improved Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (ICEEMDAN) to decompose the prediction error sequence and applies five different models to correct these errors [101].

### v2

@@ 4 For Engineering: Factors Influencing Carbon Credit Prices
In contrast, China’s carbon market is influenced by more policy and structural factors. Carbon regulation intensity is a critical factor in China’s carbon market, as the implementation of carbon regulation policies has a direct short-term impact on carbon prices, particularly by increasing compliance costs for organisations and suppressing market demand [n4]. Crude oil and coal prices also play a crucial role in China’s market, with coal price fluctuations being especially prominent, reflecting China’s reliance on coal in its energy structure [n11]. In addition, exchange rates and interest rates significantly impact China’s carbon market, particularly in the short and medium term, as fluctuations in exchange rates and interest rates affect market liquidity and energy prices [n11, n4]. Lastly, the growth of industrial production and renewable energy also has a significant influence on China’s carbon prices. These factors reflect the level of industrialization in China and its ongoing energy transition, with increases in industrial production and renewable energy generation being closely linked to carbon market prices [n7, n16].
@@ 5 Algorithms for Predicting Carbon Credit Prices
After thoroughly reviewing the factors influencing carbon prices, it becomes clear that these factors are not only critical for understanding carbon market fluctuations but also serve as essential input variables for carbon price prediction algorithms. Effectively incorporating these factors into predictive models can significantly enhance the accuracy of carbon price forecasts, making the review of these algorithms even more meaningful and necessary. By accurately predicting carbon prices, organisations can strategically plan their carbon credit purchases when prices are lower, thereby better managing their costs. This not only helps organisations control overall expenditures and strengthen their competitiveness in the market but also encourages them to actively participate in carbon management and transparently disclose their carbon emissions. In doing so, organisations can gain greater trust and advantage in markets that increasingly prioritize ESG principles. The following sections will review the latest and effective carbon price prediction algorithms.
@@
Sun and Wang used a combined method of empirical mode decomposition [38emd] and least squares support vector machine to predict carbon prices. They proposed a hybrid model based on factor analysis, empirical mode decomposition, improved particle swarm optimization [38swamoptim], and least squares support vector machine [LSVR38] to consider historical carbon prices and external factors influencing carbon prices. The study investigated three typical carbon markets in China, and the results showed that this hybrid model was more accurate in predicting carbon prices than other models [carbon38].
@@
Nadirgil employed the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) [CEEMDAN39] technique to decompose the carbon price time series into multiple Intrinsic Mode Functions (IMFs), capturing various frequency characteristics within the data. Subsequently, Permutation Entropy (PE) was utilized to classify the IMFs, grouping IMFs with similar complexity into Cooperative Intrinsic Mode Functions (Co-IMFs) to reduce model computational complexity and the risk of overfitting. For high-frequency Co-IMFs, Variational Mode Decomposition (VMD) was further applied for secondary decomposition to mitigate the impact of high-frequency noise on model prediction. Ultimately, the study integrated multiple machine learning models, including Long Short-Term Memory (LSTM), Gated Recurrent Unit (GRU), Multilayer Perceptron (MLP), and Backpropagation Neural Network (BPNN), and optimized model parameters using Genetic Algorithm (GA) to construct 48 different hybrid prediction models. The effectiveness of the dual decomposition method (CEEMDAN-VMD-BPNN-GA) in addressing prediction complexity was ultimately demonstrated [carbon39].
@@
Hao and Tian proposed an innovative carbon price prediction framework. First, the framework applies ICEEMDAN [ICEEMDAN40] to decompose the carbon price time series, effectively removing noise while retaining key signal features. Then, feature selection is conducted using the Partial Autocorrelation Function (PACF) and Maximum Relevance Minimum Redundancy (mRMR) methods. Next, the parameters of the KELM model are optimized using the Multi-Objective Chaotic Sine-Cosine Algorithm (MOCSCA). Finally, a multi-step prediction strategy is designed using the recursive method, which not only accurately predicts short-term price fluctuations but also provides reliable support for long-term trend forecasting [carbon40].
@@
Li and Liu proposed a carbon price prediction model based on secondary decomposition and feature selection. First, the carbon price time series is subjected to multi-level decomposition using Discrete Wavelet Transform (DWT) [DWT41] and ICEEMDAN, breaking it down into sub-sequences of different frequencies. Simultaneously, a three-stage feature selection method is employed to identify external influencing factors that are highly correlated with carbon prices. Subsequently, Support Vector Regression (SVR) [SVR41] and Multilayer Perceptron are used to predict the high-frequency and low-frequency sub-sequences, respectively, in order to capture short-term fluctuations and long-term trends. Finally, a weighted ensemble of the predictions from each sub-sequence is performed to obtain the overall carbon price prediction [carbon41].
@@
Wang et al. proposed a comprehensive carbon price prediction framework. Initially, the carbon price time series undergoes multi-level decomposition and denoising using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise and Wavelet Transform (WT) to extract key features and reduce noise interference. Subsequently, Principal Component Analysis (PCA), Random Forest (RF), and Gradient Boosting Decision Tree (GBDT) are employed for feature selection, ensuring that the variables input into the model possess the strongest predictive capabilities. In the prediction phase, Bidirectional Gated Recurrent Unit (BIGRU) [BIGRU42] is utilized for point prediction, leveraging bidirectional layers to capture the complex dynamic patterns within the time series. Following this, Gaussian Process Regression (GPR) [GPR42] is applied for interval prediction, providing confidence intervals for the forecast results [carbon42].
@@
Liu and Shen proposed a carbon price prediction model based on Empirical Wavelet Transform (EWT) [EWT43] and Gated Recurrent Unit, aiming to enhance prediction accuracy by integrating multiple techniques. First, the carbon price time series is decomposed using EWT, extracting components of different frequencies, including trend components, low-frequency components, and high-frequency components. Subsequently, Fuzzy C-Means (FCM) [FCM43] clustering is employed to classify these components, further refining the signal characteristics. Then, the Partial Autocorrelation Function is used to analyze the lag relationships of each component, determining the optimal lag order as input for the GRU model. Finally, the GRU model performs single-step predictions for each component, and the predicted results are aggregated to obtain the overall carbon price prediction [carbon43].
@@
Huang and He proposed a carbon price prediction model based on multi-scale joint prediction, integrating techniques such as Mean Optimization Empirical Mode Decomposition (MOEMD), Grey Relational Analysis (GRA), Factor Analysis (FA), Cooperative Kidney Algorithm (CKA) [cka44], and Extreme Learning Machine (ELM). First, MOEMD is used to decompose the carbon price data, extracting modal components at different frequencies. Then, GRA and FA are employed to select both structured and unstructured data highly correlated with carbon price fluctuations, including unstructured data like Baidu Search Index. Subsequently, CKA is utilized to optimize the parameters of the ELM model, enhancing prediction accuracy and stability. The experimental results demonstrate that the proposed MOEMD-CKA-ELM model significantly outperforms other comparative models in carbon price prediction [carbon44].
@@
Yang et al. proposed a carbon price prediction model based on a multi-factor combination, integrating Modified Ensemble Empirical Mode Decomposition (MEEMD), Long Short-Term Memory networks [lstm45], and a rule-based machine reasoning system. First, MEEMD is employed to decompose the carbon price time series, extracting modal components of different frequencies, which provide clear signals for subsequent predictions [meemd46]. Then, the LSTM model, with its strong capability for processing time series data, accurately predicts each modal component. To further enhance prediction accuracy, a rule-based machine reasoning system is introduced to automatically optimize the key parameters of the LSTM model, ensuring robust performance under various conditions. The experimental results demonstrate that the proposed MEEMD-LSTM model outperforms traditional models in carbon price prediction [carbon45].
@@
Lu et al. proposed a carbon price and trading volume prediction framework based on multi-model comparison, incorporating the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise technique to evaluate the performance of six different machine learning models. The machine learning models used include XGBoost [xgboost47], Random Forest[random48], Grey Wolf Optimizer-based Nonlinear Extension of the Arps Decline Model (GWO-KNEA) [knea49], Particle Swarm Optimization-based Support Vector Machine (PSO-SVM) [psosvm50], Simulated Annealing and Fruit Fly Optimization Algorithm-based Support Vector Machine (SA-FFOA-SVM), and Radial Basis Function Neural Network (RBFNN) [rbfnn51]. Through experiments conducted on data from eight carbon trading markets in China, the authors compared the prediction accuracy and stability of each model. The study found that the CEEMDAN-GWO-KNEA and CEEMDAN-RBFNN models performed exceptionally well across multiple datasets [carbon47].
@@
Tian et al. proposed a novel hybrid model based on feature selection and multi-objective optimization algorithms for carbon price prediction. The study used Weighted Regularized Extreme Learning Machine (WRELM) as the artificial intelligence algorithm to improve the accuracy and stability of carbon price prediction [wrema52]. The model overcame the limitations of statistical models in handling carbon price prediction and optimized the model using the Multi-Objective Grasshopper Optimization Algorithm (MOGOA), achieving excellent prediction results [mogoa53]. This research provides important references for the rational and stable development of the carbon market [carbon48].
@@
Yang et al. proposed a carbon price prediction model based on Modified Ensemble Empirical Mode Decomposition, Long Short-Term Memory networks, and an Improved Whale Optimization Algorithm (IWOA) [iwoa56]. First, MEEMD is used to decompose the carbon price time series into several Intrinsic Mode Functions and a residual component, reducing noise and mode mixing issues [meemd54]. Subsequently, the LSTM model is employed to predict the future values of each IMF and the residual component, capturing long-term dependencies in the time series [lstm45]. To further enhance prediction performance, IWOA is used to optimize the initial weights and biases of the LSTM model [carbon49].
@@
Zhao et al. proposed a carbon price prediction model that combines the Hodrick-Prescott (HP) filter, an improved Grey Model (Variable Grey Model, VGM), and Extreme Learning Machine [vgm57]. First, the HP filter is used to decompose the carbon price data and its influencing factors into long-term trends and short-term fluctuations, which are modeled separately. The VGM model is then employed to predict the long-term trends, while the ELM handles the short-term fluctuations. By combining the predictions from these two components, the model effectively enhances the accuracy of carbon price prediction [carbon51].
@@
Wang et al. proposed a carbon price interval prediction model based on Lower Upper Bound Estimation (LUBE), Asymmetric Multi-Objective Evolutionary Algorithm (MOEA), and Long Short-Term Memory networks [lstm45]. First, the carbon price time series is decomposed using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise to reduce noise and extract key features. Next, Symbol Transfer Entropy (STE) is employed for causal inference and feature selection to ensure the effectiveness of the input variables [ste58]. Then, based on the prediction points generated by LSTM, the LUBE model uses MOEA to optimize parameters and generate asymmetric prediction intervals [np17].
@@
Deng et al. proposed a carbon price trend prediction model that integrates several advanced techniques. The model employs Complete Ensemble Empirical Mode Decomposition with Adaptive Noise to decompose carbon price time series into Intrinsic Mode Functions and residuals, allowing the extraction of significant features. Following this, a Two-Stage Feature Selection (TFS) method is used to identify the most relevant features from these IMFs, which are then fed into a Light Gradient Boosting Machine (LightGBM) model [lightgbm60]. The hyperparameters of the LightGBM model are optimized using Bayesian Optimization Algorithm (BOA) to enhance its predictive performance [boa61]. Finally, SHapley Additive exPlanations (SHAP) is applied to interpret the model’s predictions [shap62], providing insights into the contributions of various features to the predicted carbon price trends [np16].
@@
Shi et al. proposed a deep learning framework combining Convolutional Neural Networks (CNN) and Long Short-Term Memory networks to predict carbon prices with high accuracy and robustness. The model first uses CNN layers to extract spatial features from the carbon price time series data, which are then passed to LSTM layers to capture temporal dependencies [cnn63]. This hybrid CNN-LSTM model is designed to address the limitations of traditional models in capturing complex nonlinear patterns in carbon markets [lstm64]. The study demonstrates that the CNN-LSTM model significantly outperforms single CNN or LSTM models in both prediction accuracy and robustness, particularly in handling anomalies and outliers in the data [np15].
@@
Zhang et al. proposed a hybrid model for multi-step carbon price prediction, integrating Extremely Randomized Trees (ET), Multivariate Variational Mode Decomposition (MVMD), and Long Short-Term Memory networks. The model employs ET to select the most influential external factors affecting carbon price fluctuations, simplifying the model’s complexity and enhancing prediction accuracy. MVMD [mvmd65] is then used to decompose the carbon price and selected external variables into multiple modes, allowing for the extraction of stable frequency components that are crucial for accurate predictions. Each mode is subsequently modeled using LSTM to capture long-term dependencies in the time series, with the final prediction obtained by aggregating the individual predictions [np14].
@@
Zhang et al. proposed an enhanced decomposition-ensemble model for both point and probabilistic carbon price forecasting. The model integrates Complete Ensemble Empirical Mode Decomposition with Adaptive Noise, Fuzzy Entropy (FE) [fe67], Variational Mode Decomposition [vmd66], Long Short-Term Memory networks, Gated Recurrent Unit networks, Cuckoo Search Algorithm (CS) [cs68], and Gaussian Process Regression. Initially, CEEMDAN is employed to decompose the carbon price series into Intrinsic Mode Functions and residuals, followed by FE to quantify the complexity of these IMFs. High-complexity IMFs undergo further decomposition using VMD. LSTM and GRU models then predict each decomposed component, with CS optimizing their weights. Finally, GPR is used to generate prediction intervals, offering probabilistic insights [np13].
@@
Hong et al. proposed a hybrid carbon price forecasting framework that integrates a Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer) with a Multimodel Optimization Piecewise Error Correction System [da-fed69]. The DA-FEDformer model enhances frequency domain information extraction using a multi-layer perceptron integrated into the encoder-decoder layers, combined with an Improved Kernel Mean Square Error (IKMSE) loss function and the Lion Optimizer. This model generates initial carbon price predictions, which are then refined by the error correction system [ICEEMDAN40]. The system employs Improved Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (ICEEMDAN) to decompose the prediction error sequence and applies five different models to correct these errors [np2].
Xu et al. proposed a hybrid framework for real-time carbon price prediction, tackling data feature drift and noise challenges in time series decomposition. The model uses Real-Time Complete Ensemble Empirical Mode Decomposition (RT-EMD) to decompose time series into meaningful components and residuals, which are further processed with Fuzzy Information Granulation (FIG) to reduce complexity. Key features are selected using Mutual Information Association Features (MIAF), and predictions are made using a Regularized Extreme Learning Machine (RELM) for deterministic forecasting and Gaussian Process Regression for probabilistic intervals [xu2024novel].
Zheng et al. proposed a decomposition-ensemble model for carbon price point and interval forecasting, integrating Sequential Variational Mode Decomposition (SVMD), Sample Entropy (SE), Singular Spectrum Analysis (SSA), and machine learning models. Initially, SVMD is employed to decompose the carbon price time series into Intrinsic Mode Functions and residuals. Subsequently, SE is utilized to evaluate the complexity of these IMFs, classifying them into low-frequency and high-frequency components. During the point forecasting stage, low-frequency components are predicted using a CatBoost model optimized with SSA, while high-frequency components are predicted using a Kernel Extreme Learning Machine, also optimized with SSA. After completing the predictions for both low-frequency and high-frequency components, their results are combined through direct summation to generate the final point forecasts. Finally, Adaptive Bandwidth Kernel Density Estimation (ABKDE) is applied to the residuals to estimate the error distribution and generate probabilistic intervals, providing uncertainty quantification and robust forecasting results [zheng2024multifactor].
Mao and Yu proposed a hybrid carbon price forecasting model integrating CEEMDAN, VMD, Sample Entropy, LightGBM, and Grey Wolf Optimizer (GWO). The model employs CEEMDAN to decompose the carbon price series into Intrinsic Mode Functions and residuals, followed by VMD for further refinement of high-noise IMFs. SE is then used to classify the IMFs into high-frequency, medium-frequency, and low-frequency components, which are individually predicted using LightGBM models optimized by GWO. The predictions from all components are aggregated into the final point forecast, with SHAP applied to ensure model interpretability [mao2024hybrid].
Zhang et al. proposed a text-based carbon price forecasting framework based on a multivariate temporal graph neural network (MtemGNN), aiming to enhance prediction accuracy by integrating unstructured textual information from online news headlines [mtemgnn2025]. First, the framework incorporates historical carbon price data together with online news headline data as multi-source inputs. To extract effective textual features, a Dynamic Topic Model (DTM) is employed to capture time-varying thematic information from news across different periods, while sentiment information is extracted using SnowNLP and further optimized by introducing an accumulative decay factor to account for the temporal attenuation of news impact. Subsequently, a multivariate temporal graph neural network is constructed to automatically learn the complex interdependencies among carbon prices, news topics, and sentiment features in a data-driven manner without requiring prior knowledge of graph topology. By simultaneously modeling temporal autocorrelation and inter-feature relationships, the proposed MtemGNN framework demonstrates superior carbon price forecasting performance compared with multiple benchmark models across several regional carbon markets in China [mtemgnn2025].
Ji et al. proposed a novel probabilistic carbon price prediction framework by integrating mixed-frequency modeling with a Transformer-based quantile regression architecture, aiming to improve both point forecasting accuracy and uncertainty quantification of carbon prices [qrtransformer_midas2025]. First, the framework incorporates multi-source explanatory variables with heterogeneous sampling frequencies, including energy, financial, macroeconomic, and environmental factors, which are transformed into unified representations through a mixed-data sampling (MIDAS) approach. This mixed-frequency modeling strategy effectively preserves high-frequency information while maintaining consistency with low-frequency carbon price series. Subsequently, a Transformer network is employed to capture complex nonlinear temporal dependencies within the mixed-frequency features, leveraging its self-attention mechanism to model long-range interactions in the carbon price dynamics. To further characterize predictive uncertainty, quantile regression is embedded into the Transformer framework, enabling the direct estimation of carbon price distributions at different quartiles rather than relying solely on point predictions. Finally, kernel density estimation is applied to the quantile-based outputs to construct continuous probabilistic forecasts. Empirical results demonstrate that the proposed MIDAS–Transformer–quantile framework outperforms multiple benchmark models in terms of both predictive accuracy and interval reliability across several carbon markets [qrtransformer_midas2025].
Wang et al. proposed a carbon trading price prediction framework based on a dual decomposition strategy that integrates Complex Wavelet Transform (CWT) and Gaussian Process Regression (GPR), aiming to enhance prediction accuracy by effectively handling nonlinear and non-stationary characteristics of carbon price series [cwtgpr_twicedec_2025]. First, the carbon price time series is decomposed using CWT to extract multi-scale components, with Kolmogorov complexity employed to determine the optimal number of decomposition layers and Approximate Entropy used to reconstruct high-frequency components with significant randomness. Subsequently, the residual sequences obtained from the first decomposition are further decomposed through GPR, which separates trend and high-frequency kernel components by maximizing the log marginal likelihood, forming a second-stage decomposition process. After the twice decomposition, a Gated Recurrent Unit (GRU) model is utilized to predict each decomposed component, while the Crested Porcupine Optimization (CPO) algorithm is applied to optimize the model parameters. Finally, the predicted results of all components are aggregated to generate the final carbon trading price forecasts. Experimental results based on carbon trading data from Wuhan and Beijing demonstrate that the proposed CWT–GPR–GRU framework achieves superior predictive performance compared with several benchmark models, particularly in short-term forecasting horizons [cwtgpr_twicedec_2025].